OpenStudy (anonymous):

help 4 medals!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! @iGreen

OpenStudy (anonymous):

Match each shape with its surface area. Use three point one four for pi and round to the nearest whole number where necessary. Match Term Definition square pyramid: base length ten meters, slant height 6 meters A) three hundred one square meters cone: radius 8 meters, slant height 4 meters B) two hundred thirty-one square meters square pyramid: base length eleven meters, slant height 5 meters C) two hundred twenty square meters cone: radius 5 meters, slant height 7 meters D) one hundred eighty-eight square meters

OpenStudy (igreen):

Surface Area of Square Pyramid: \(SA = b^2 + 2 × b × s\) Where b is base length r is radius, and s is slant height. Surface Area of Cone: \(SA = \pi r s + \pi r^2\) Where r is radius, and where s is slant height. Just plug the values into these formulas, then solve.

OpenStudy (anonymous):

wha?

OpenStudy (igreen):

If you see those question marks, reload the page.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

\[\LARGE\]\(THANKU\)

OpenStudy (igreen):

Np.